# Ramification filtration via deformations, II

### v3

@@ 2.6. The action $\Omega_{U}$
The derivation $B^{{\dagger}}$ can be recovered via methods from [9]-[10].
@@
where $A^{\,{\dagger}}_{U}=\widetilde{\exp}(UB^{\,{\dagger}})$. As a matter of fact, when obtaining (2.2) we are working inside $\oplus_{m\geqslant 0}U^{j}\widetilde{\mathcal{N}}^{{\dagger}}[j,m]$.
It is easy to see that we have action of the additive group $\mathbb{G}_{a}=\mathrm{Spec}\,\mathbb{Z}/p^{M}[U]$ on $\widetilde{\mathcal{N}}^{\,{\dagger}}$. Similarly to [12] and by taking into account the compatibility of this action with the double grading $\widetilde{\mathcal{N}}^{\,{\dagger}}[j,m]$ we obtain the following:
@@
Prop. 2.9 is directly related to Corollary 2.8 as follows:
@@ 2.7. Ideals ¯$\bar{\mathcal{L}}^{\,{\dagger}[v_{0}]}$ and ¯$\bar{\mathcal{L}}^{[v_{0}]}$
Recall that ¯$\bar{\mathcal{L}}^{\,{\dagger}}[0]$ is the minimal Lie subalgebra of $\bar{\mathcal{L}}^{\,{\dagger}}$ containing all $D^{\,{\dagger}}_{\iota n}$ with $\mathrm{ch}(\iota)=0$. By Remark from Sect. 2.3, $\mathcal{L}^{{\dagger}}[0]_{W_{M}(k)}$ is generated by
@@
Recall also that $\mathcal{L}^{{\dagger}}[0]$ is naturally identified with the Lie algebra $\mathcal{L}$ via $j^{0}:D^{{\dagger}}_{qp^{M-1}a,n}\mapsto D_{an}$. This identification agrees with filtrations and plays a crucial role below, where we work with criterion from Prop. 3.2 in the algebra $\mathcal{L}^{{\dagger}}[0]$.
@@
For each $j\geqslant 1$ set $\bar{\mathcal{N}}^{{\dagger}}\langle j\rangle=\oplus_{m}\bar{\mathcal{N}}^{{\dagger}}[j,m]$, and let $\pi:\bar{\mathcal{N}}\to\widetilde{\mathcal{N}}^{{\dagger}}$ be a natural projection.
@@
There are unique elements $\bar{c}^{(j)}\in\bar{\mathcal{N}}^{{\dagger}}\langle j\rangle$ such that $\pi(\bar{c}^{(j)})=\widetilde{c}^{(j)}$, cf. Prop. 2.9, and
$\Omega(\bar{e}^{\,{\dagger}})=\sigma(\bar{c})\circ(A^{{\dagger}}\otimes\mathrm{id}_{O^{\prime}_{M}(\mathcal{K}^{\prime})}\bar{e}^{\,{\dagger}}\circ(-\bar{c})\,$
@@
It follows easily from the fact that $\mathrm{Ker}\,\pi\subset\bar{\mathcal{L}}^{{\dagger}}_{\mathrm{m}^{\prime}}$ and for any $j$ and $m$, the map $\sigma-\mathrm{id}$ is a bijection on $\mathrm{Ker}\pi\cap\bar{\mathcal{N}}^{{\dagger}}[j,m]$. ∎
@@ 3.1. Groups and Lie algebras of nilpotent class $<p$, [24]
Let $L$ be a finite Lie $\mathbb{Z}/p^{M}$-algebra of nilpotent class $<p$, i.e. $C_{p}(L)=0$. Let $A$ be an enveloping algebra of $L$. There is a natural embedding $L\subset A/J^{p}$ and in terms of this embedding the Campbell-Hausdorff formula appears as
@@
where $\widetilde{\exp}(l_{1})\cdot\widetilde{\exp}(l_{2})\equiv\widetilde{\exp}(l_{1}\circ l_{2})\,\mathrm{mod}\,J^{p}$. This composition law provides the set $L$ with a group structure and we denote this group by $G(L)$. Clearly, $G(L)$ has period $p^{M}$ and nilpotent class $<p$. Then the correspondence $L\mapsto G(L)$ induces equivalence of the category of finite $p$-groups of period $p^{M}$ and nilpotent class $s<p$ and the category of finite Lie $\mathbb{Z}/p^{M}$-algebras of the same nilpotent class $s$. This equivalence can be naturally extended to similar categories of pro-finite ubjects.
@@ 3.2. Epimorphism $\eta^{0}$
Let $L$ be a finite $\mathbb{Z}/p^{M}$-Lie algebra of nilpotent class $<p$. Set $L_{sep}:=L_{O_{M}(\mathcal{K}_{sep})}$. Then the elements of $\mathcal{G}=\mathrm{Gal}(\mathcal{K}_{sep}/\mathcal{K})$ and $\sigma$ act on $L_{sep}$ through the second factor, $L_{sep}|_{\sigma=\mathrm{id}}=L$ and $(L_{sep})^{\mathcal{G}}=L_{\mathcal{K}}$. The covariant nilpotent Artin-Schreier theory states that for any $e\in G(L_{\mathcal{K}})$, the set
@@
Suppose $\mathcal{L}$ is the Lie algebra from Section 1.1 and $e=e(t)\in\mathcal{L}_{\mathcal{K}}$ satisfies $e\equiv\sum_{a\in\mathbb{Z}^{0}(p)}t^{-a}D_{a0}\,\mathrm{mod}\,C_{2}(\mathcal{L})_{O_{M}(\mathcal{K})}$. Recall that for $n\in\mathbb{Z}/N_{0}$, $D_{0n}$ is defined as $\sigma^{n}(\alpha_{0})D_{0}$, where $\alpha_{0}=\hat{\alpha}_{0}\,\mathrm{mod}p^{M}$ and $\alpha_{0}\in W(k)$ has absolute trace 1. (We followed this agreement in [1]-[3] to relate the identification $\eta^{0}\,\mathrm{mod}\,C_{2}(\mathcal{G})$, cf. below, to the identification of class field theory.) Then a profinite version of the above Artin-Schreier theory gives the existence of $f\in\mathcal{L}_{sep}$ such that $\sigma f=e\circ f$ and the map $\eta^{0}_{f}:=\pi_{f}(e)\,\mathrm{mod}\,\mathcal{G}^{p^{M}}C_{p}(\mathcal{G})$ induces a group isomorphism $\eta^{0}_{f}:\mathcal{G}_{<p}\simeq G(\mathcal{L})$. We fix a choice of such $f$ and use the simpler notation $\eta^{0}=\eta^{0}_{f}$.
@@
$\bf A)$ Let $\mathcal{N}$ be a $W_{M}(k)$-Lie subalgera of $\mathcal{L}_{O_{M}(\mathcal{K})}$ generated by all $t^{-a}D_{an}$, $a\in\mathbb{Z}^{0}(p)$. Then $e\in\mathcal{N}$ and $e\equiv\sum_{a\in\mathbb{Z}^{0}(p)}t^{-a}D_{a0}\,\mathrm{mod}\,[\mathcal{N},\mathcal{N}]$.
@@
The above quite general result will be explicitly specified in Sec. 4, where we present another proof of the explicit description of $\mathcal{L}^{[v_{0}]}$ in terms of generators $\{D_{an}\ |\ a\in\mathbb{Z}^{+}(p),n\in\mathbb{Z}/N_{0}\}\cup\{D_{0}\}$ of $\mathcal{L}_{W_{M}(k)}$ from [3].
@@ 3.3. The criterion
Consider the parameters $N^{*},r^{*},q$ from Sec. 2.1.
Consider the field etension $\mathcal{K}^{\prime}/\mathcal{K}$ from Sec. 2.4. Recall $\mathcal{K}^{\prime}=k((t_{01}))$ and the local parameter $t_{01}$ was used to construct the lifts $\ O^{\prime}_{M}(\mathcal{K}^{\prime})$ and $O^{\prime}_{M}(\mathcal{K}_{sep})$. There is a natural embedding $\iota:\sigma^{M-1}O_{M}(\mathcal{K})\to O^{\prime}_{M}(\mathcal{K}^{\prime})$. This allows us to identify the Lie $O^{\prime}_{M}(\mathcal{K}^{\prime})$-algebra $\mathcal{N}^{{\dagger}}[0]$ with subalgebra of $(\mathcal{L}_{\sigma^{M-1}O_{M}(\mathcal{K})})_{O^{\prime}_{M}(\mathcal{K}^{\prime})}$. In particular, $\sigma^{M-1}e(t)$ is identified with $\Omega\left(\sigma^{M-1}e(t_{1}^{q})\right)$.
@@
For $Y\in\mathcal{L}_{O^{\prime}_{M}(\mathcal{K}_{sep})}$ and an ideal $\mathcal{I}$ in $\mathcal{L}$, define the field of definition of $Y\,\mathrm{mod}\,\mathcal{I}_{O^{\prime}_{M}(\mathcal{K}_{sep})}$ over $\mathcal{K}^{\prime}$ as $\mathcal{K}^{\prime}(Y\,\mathrm{mod}\,\mathcal{I}_{O^{\prime}_{M}(\mathcal{K}_{sep})}):=\mathcal{K}_{sep}^{\mathcal{H}}$, where =$\mathcal{H}=\{g\in\mathcal{G}^{\prime}=\mathrm{Gal}(\mathcal{K}_{sep}/\mathcal{K}^{\prime})\ |\ (\mathrm{id}_{\mathcal{L}}\otimes g)Y\equiv Y\,\mathrm{mod}\,\mathcal{I}_{sep}\}$.
For a field extension $\mathcal{E}^{\prime}/\mathcal{K}^{\prime}$ in $\mathcal{K}_{sep}$, define the biggest upper ramification number
$v(\mathcal{E}^{\prime}/\mathcal{K}^{\prime})=\mathrm{max}\{v\ |\ \mathrm{Gal}(\mathcal{K}_{sep}/\mathcal{K}^{\prime})^{(v)}\text{acts non-trivially on }\mathcal{E}^{\prime}\}\,.$
The methods from [1, 2, 3] are based on the following criterion.
Set $e_{1}^{(q)}:=e(t_{1}^{q})$. Then $\Omega\left(\sigma^{M-1}e_{1}^{(q)}\right)=e(t_{1}^{qp^{M-1}}E(t_{1}^{b^{*}})^{-p^{M-1}})$.
@@
The above inequality can be also rewritten in the following way $v(\mathcal{K}^{\prime}(X\,\mathrm{mod}\,\mathcal{I}_{O^{\prime}_{M}(\mathcal{K}_{sep})})/\mathcal{K})<v_{0}\,.$
Suppose $\widetilde{\mathcal{J}}\subset\mathcal{L}$ is a closed ideal and $\pi:\mathcal{L}\longrightarrow L:=\mathcal{L}/\widetilde{\mathcal{J}}$ is a natural projection. Then we can use the appropriate images $e_{L},e_{L}^{(q)}\in L_{O^{\prime}_{M}(\mathcal{K}^{\prime})}$ and $f_{L},X_{L}\in L_{O^{\prime}_{M}(\mathcal{K}_{sep})}$ to state the following analog of Prop.3.2 for the ideal $L^{(v_{0})}:=\pi(\mathcal{L}^{(v_{0})})$ of $L$.
@@ 3.4. Relation between $\mathcal{L}^{(v_{0})}$ and $\mathcal{L}^{[v_{0}]}$
Since $\mathcal{L}^{{\dagger}}[0](p)=\mathrm{Ker}(\mathcal{L}^{{\dagger}}[0]\to\bar{\mathcal{L}}^{{\dagger}}[0])\subset\mathcal{L}^{[v_{0}]}$ we can take any $x\in\mathcal{N}^{{\dagger}}[0]$ such that $\bar{x}=x\,\mathrm{mod}\,\mathcal{L}^{{\dagger}}[0](p)_{O^{\prime}_{M}(\mathcal{K}^{\prime})}$ and consider congruence (3.3) in $\mathcal{N}^{{\dagger}}[0]$:
@@
The following proposition will imply the statement of Theorem 3.1 when $w=p$.
@@
Use induction on $w\geqslant 1$. If $w=1$ there is nothing to prove. So, assume that proposition holds for some $1\leqslant w<p$.
For non-commuting variables $U$ and $V$, let
@@
Use that $\Omega\left(\sigma^{M-1}e_{1}^{(q)}\right)\circ X=\sigma X\circ\sigma^{M-1}e_{1}^{(q)}$ and there are the following congruences modulo $\left([\mathcal{L}^{(v_{0})},\mathcal{L}]+C_{w_{0}+1}(\mathcal{L})\right)_{sep}$:
@@
In order to prove part $b_{w+1})$ we need to verify the following condition:
$(\bullet_{1})$ if $qa\geqslant b^{*}U(w,m)$ then $p^{m}D_{a0}\in(\mathcal{L}^{(v_{0})}+C_{w+1}(\mathcal{L}))_{W_{M}(k)}$.
This will be done by the use of the condition
(3.4) $v(\mathcal{K}^{\prime}(X\,\mathrm{mod}\,\mathcal{L}^{(v_{0})}_{sep})/\mathcal{K}^{\prime})<qv_{0}-b^{*}\,.$
Suppose $y=\sum_{\iota}y_{\iota}+y_{+}$, where $y_{+}\in\mathcal{L}_{\mathrm{m}^{\prime}}$, $\iota$ runs over $\mathfrak{A}^{+}(p)$ and the corresponding $y_{\iota}$ appear in the form $y_{\iota}=\sum_{r\geqslant 0}t_{1}^{-\iota p^{r}}l_{br}$ with all $l_{\iota r}\in\mathcal{L}_{W_{M}(k)}$.
Then $X=\sum_{\iota}X_{\iota}+X_{+}$ where $\sigma X_{\iota}-X_{\iota}=y_{\iota}$ and $\sigma X_{+}-X_{+}=y_{+}$. Here $X_{+}\in\mathcal{L}_{\mathrm{m}^{\prime}}$ and does not contribute to (3.4).
The elements $X_{\iota}$ contribute to (3.4) as follows:
if $\iota\in\mathfrak{A}^{+}(p)$ and $\iota/p^{v_{p}(\iota)}\geqslant qv_{0}-b^{*}$ then
$\sum_{r}\sigma^{-r}l_{\iota r}\in(\mathcal{L}^{(v_{0})}+C_{w+1}(\mathcal{L}))_{W_{M}(k)}\,.$
We can follow the contribution to upper ramification in (3.4) for each $\iota\in\mathfrak{A}^{+}(p)$ separately because by Proposiion 2.4 these contributions are independent.
The element $y\in\mathcal{N}^{{\dagger}}[0]\{1\}$ has a linear part
(3.5) $L(y):=\Omega\left(\sigma^{M-1}e_{1}^{(q)}\right)-\sigma^{M-1}e_{1}^{(q)}=\sum_{a,s}t_{1}^{-ap^{M-1}q}aS_{b^{*}}^{s}D_{a0}$
and a non-linear part, which is a $W_{M}(k)$-linear combination of various

### v5

@@ 2.6. The action $\Omega_{U}$
The derivation $B^{{\dagger}}$ can be recovered via methods from papers [9]-[10].
@@
where $A^{\,{\dagger}}_{U}=\widetilde{\exp}(UB^{\,{\dagger}})$.
It is easy to see that we have the action of the additive group $\mathbb{G}_{a}=\mathrm{Spec}\,\mathbb{Z}/p^{M}[U]$ on $\widetilde{\mathcal{N}}^{\,{\dagger}}$. Similarly to [12] and by taking into account the compatibility of this action with the grading on $\widetilde{\mathcal{N}}^{\,{\dagger}}$ we obtain the following:
@@
Prop. 2.11 is directly related to Corollary 2.10 as follows:
@@ 2.7. Ideals ¯$\bar{\mathcal{L}}^{\,{\dagger}[v_{0}]}$ and ¯$\bar{\mathcal{L}}^{[v_{0}]}$
Recall that ¯$\bar{\mathcal{L}}^{\,{\dagger}}[0]$ is the minimal Lie subalgebra of $\bar{\mathcal{L}}^{\,{\dagger}}$ containing all $D^{\,{\dagger}}_{\iota n}$ with $\mathrm{ch}(\iota)=0$. By Remark from Sect. 2.3, $\mathcal{L}^{{\dagger}}[0]_{W(k)}$ is generated by
@@
Recall also that $\mathcal{L}^{{\dagger}}[0]$ is naturally identified with a Lie subalgebra of $\mathcal{L}$ via $j^{0}:D^{{\dagger}}_{qp^{M-1}a,n}\mapsto D_{an}$. This identification agrees with filtrations and plays a crucial role below, where we work with the criterion from Prop. 3.2 in the algebra $\mathcal{L}^{{\dagger}}[0]$.
@@
There are unique elements $\bar{c}^{(j)}\in\bar{\mathcal{N}}^{{\dagger}}(j)$ such that $\pi(\bar{c}^{(j)})=\widetilde{c}^{(j)}$ and
$\Omega(\bar{e}^{\,{\dagger}})=\sigma(\bar{c})\circ(A^{{\dagger}}\otimes\mathrm{id}_{O^{\prime}_{M}(\mathcal{K}^{\prime})})\bar{e}^{\,{\dagger}}\circ(-\bar{c})\,$
@@
It follows easily from the fact that $\bar{\mathcal{N}}^{{\dagger}}(p-1)\subset\bar{\mathcal{L}}^{{\dagger}}_{\mathrm{m}^{\prime}}$ and, therefore, the map $\sigma-\mathrm{id}$ is a bijection on $\bar{\mathcal{N}}^{{\dagger}}(p-1)$. ∎
@@ 3. Relation to Galois theory
Recall that $\mathcal{K}=k((t_{0}))$ and we use the construction of the lift $O_{M}(\mathcal{K}_{sep})$ of $\mathcal{K}_{sep}$ modulo $p^{M}$ with respect to our given uniformiser $t_{0}$. This lift appears in the form $W_{M}(\sigma^{M-1}\mathcal{K}_{sep})[t]\subset W_{M}(\mathcal{K}_{sep})$, where $t=[t_{0}]$ is the Teichmuller representative of $t_{0}$. Then $\mathcal{G}=\mathrm{Gal}(\mathcal{K}_{sep}/\mathcal{K})$ acts on this lift and $O_{M}(\mathcal{K}_{sep})^{\mathcal{G}}=O_{M}(\mathcal{K})=W_{M}(k)((t))$.
@@ 3.1. Groups and Lie algebras of nilpotent class $<p$, [25]
Let $L$ be a finite Lie $\mathbb{Z}/p^{M}$-algebra of nilpotent class $<p$, i.e. $C_{p}(L)=0$. Let $A$ be an enveloping algebra of $L$. There is a natural embedding $L\subset A/J^{p}$ (here $J$ is the augmentation ideal of $A$) and in terms of this embedding the Campbell-Hausdorff formula appears as
@@
where $\widetilde{\exp}(l_{1})\cdot\widetilde{\exp}(l_{2})\equiv\widetilde{\exp}(l_{1}\circ l_{2})\,\mathrm{mod}\,J^{p}$. This composition law provides the set $L$ with a group structure and we denote this group by $G(L)$. Clearly, $G(L)$ has period $p^{M}$ and nilpotent class $<p$. Then the correspondence $L\mapsto G(L)$ induces equivalence of the category of finite $p$-groups of period $p^{M}$ and nilpotent class $s<p$ and the category of finite Lie $\mathbb{Z}/p^{M}$-algebras of the same nilpotent class $s$. This equivalence can be naturally extended to similar categories of pro-finite objects.
@@ 3.2. Epimorphism $\eta^{0}$
Let $L$ be a finite $\mathbb{Z}/p^{M}$-Lie algebra of nilpotent class $<p$. Set $L_{sep}:=L_{O_{M}(\mathcal{K}_{sep})}$. Then the elements of $\mathcal{G}=\mathrm{Gal}(\mathcal{K}_{sep}/\mathcal{K})$ and Frobenius $\sigma$ act on $L_{sep}$ through the second factor, $L_{sep}|_{\sigma=\mathrm{id}}=L$ and $(L_{sep})^{\mathcal{G}}=L_{O_{M}(\mathcal{K})}$. The covariant nilpotent Artin-Schreier theory states that for any $e\in G(L_{O_{M}(\mathcal{K})})$, the set
@@
Suppose $\mathcal{L}$ is the Lie algebra from Section 1.1 and $e=e(t)\in\mathcal{L}_{\mathcal{K}}$ satisfies $e\equiv\sum_{a\in\mathbb{Z}^{0}(p)}t^{-a}D_{a0}\,\mathrm{mod}\,C_{2}(\mathcal{L})_{O_{M}(\mathcal{K})}$. Recall that for $n\in\mathbb{Z}/N_{0}$, $D_{0n}$ is defined as $\sigma^{n}(\alpha_{0})D_{0}$, where $\alpha_{0}=\hat{\alpha}_{0}\,\mathrm{mod}\,p^{M}$ and $\hat{\alpha}_{0}\in W(k)$ has absolute trace 1. (We followed this agreement in [1]-[3] to relate the identification $\eta^{0}\,\mathrm{mod}\,C_{2}(\mathcal{G})$, cf. below, to the identification of class field theory.) Then a profinite version of the above Artin-Schreier theory gives the existence of $f\in\mathcal{L}_{sep}$ such that $\sigma f=e\circ f$ and the map $\eta^{0}_{f}:=\pi_{f}(e)\,\mathrm{mod}\,\mathcal{G}^{p^{M}}C_{p}(\mathcal{G})$ induces a group isomorphism $\eta^{0}_{f}:\mathcal{G}_{<p}\simeq G(\mathcal{L})$. We fix a choice of such $f$ and use the simpler notation $\eta^{0}=\eta^{0}_{f}$.
@@
$\bf A)$ Let $\mathcal{N}$ be a $W(k)$-Lie subalgera of $\mathcal{L}_{O_{M}(\mathcal{K})}$ generated by all $t^{-a}D_{an}$, $a\in\mathbb{Z}^{0}(p)$. Then $e\in\mathcal{N}$ and $e\equiv\sum_{a\in\mathbb{Z}^{0}(p)}t^{-a}D_{a0}\,\mathrm{mod}\,[\mathcal{N},\mathcal{N}]$.
@@
The above quite general result will be explicitly specified in Sec. 4, where we present another proof of the explicit description of $\mathcal{L}^{[v_{0}]}$ in terms of generators $\{D_{an}\ |\ a\in\mathbb{Z}^{+}(p),n\in\mathbb{Z}/N_{0}\}\cup\{D_{0}\}$ of $\mathcal{L}_{W(k)}$ from [3].
@@ 3.3. The criterion
Consider the parameters $N^{*},r^{*},q$ from Section 2.1 and the field extension $\mathcal{K}^{\prime}/\mathcal{K}$ from Section 2.4. Recall that $\mathcal{K}^{\prime}=k((t_{01}))$ and we use the local parameter $t_{01}$ to construct the lifts $\ O^{\prime}_{M}(\mathcal{K}^{\prime})=W_{M}(k)((t_{1}))$ and $O^{\prime}_{M}(\mathcal{K}^{\prime}_{sep})=W_{M}(\sigma^{M-1}\mathcal{K}^{\prime})[t_{1}]$, where $t_{1}=[t_{01}]$. There is a natural embedding of $W(k)$-algebras $\iota:\sigma^{M-1}O_{M}(\mathcal{K})\to O^{\prime}_{M}(\mathcal{K}^{\prime})$. This allows us to identify the Lie $W(k)$-algebra $\mathcal{N}^{{\dagger}}[0]$ with subalgebra of $\mathcal{L}_{O^{\prime}_{M}(\mathcal{K}^{\prime})}$. In particular, $\sigma^{M-1}e(t)$ is identified with $\Omega\left(\sigma^{M-1}e(t_{1}^{q})\right)$, cf. Section 2.5.
@@
For $Y\in\mathcal{L}_{O^{\prime}_{M}(\mathcal{K}_{sep})}$ and an ideal $\mathcal{I}$ in $\mathcal{L}$, define the field of definition of $Y\,\mathrm{mod}\,\mathcal{I}_{O^{\prime}_{M}(\mathcal{K}_{sep})}$ over $\mathcal{K}^{\prime}$ as $\mathcal{K}^{\prime}(Y\,\mathrm{mod}\,\mathcal{I}_{O^{\prime}_{M}(\mathcal{K}_{sep})}):=\mathcal{K}_{sep}^{\mathcal{H}}$, where
=$\mathcal{H}=\{g\in\mathcal{G}^{\prime}=\mathrm{Gal}(\mathcal{K}_{sep}/\mathcal{K}^{\prime})\ |\ (\mathrm{id}_{\mathcal{L}}\otimes g)Y\equiv Y\,\mathrm{mod}\,\mathcal{I}_{sep}\}\,.$
For a field extension $\mathcal{E}^{\prime}/\mathcal{K}^{\prime}$ in $\mathcal{K}_{sep}$, define the biggest upper ramification number $v(\mathcal{E}^{\prime}/\mathcal{K}^{\prime})=\mathrm{max}\{v\ |\ \mathrm{Gal}(\mathcal{K}_{sep}/\mathcal{K}^{\prime})^{(v)}\text{acts non-trivially on }\mathcal{E}^{\prime}\}\,.$
Set $e_{1}^{(q)}:=e(t_{1}^{q})$. Then $\Omega\left(\sigma^{M-1}e_{1}^{(q)}\right)=e(t_{1}^{qp^{M-1}}E(t_{1}^{b^{*}})^{-p^{M-1}})$. The methods from [1, 2, 3] are based on the following criterion.
@@
The above inequality can be rewritten also as
$v(\mathcal{K}^{\prime}(X\,\mathrm{mod}\,\mathcal{I}_{O^{\prime}_{M}(\mathcal{K}_{sep})})/\mathcal{K})<v_{0}\,.$
Suppose $\widetilde{\mathcal{J}}\subset\mathcal{L}$ is a closed ideal and $\pi:\mathcal{L}\longrightarrow L:=\mathcal{L}/\widetilde{\mathcal{J}}$ is a natural projection. Then we can use the appropriate images $e_{L},e_{L}^{(q)}\in L_{O^{\prime}_{M}(\mathcal{K}^{\prime})}$ and $f_{L},X_{L}\in L_{O^{\prime}_{M}(\mathcal{K}_{sep})}$ to state the following analog of Proposition 3.2 for the ideal $L^{(v_{0})}:=\pi(\mathcal{L}^{(v_{0})})$ of $L$.
@@ 3.4. Relation between $\mathcal{L}^{(v_{0})}$ and $\mathcal{L}^{[v_{0}]}$
Since $\mathcal{L}^{{\dagger}}[0](p)=\mathrm{Ker}(\mathcal{L}^{{\dagger}}[0]\to\bar{\mathcal{L}}^{{\dagger}}[0])\subset\mathcal{L}^{[v_{0}]}$ we can take any $x\in\mathcal{N}^{{\dagger}}[0]$ such that $\bar{x}=x\,\mathrm{mod}\,\mathcal{L}^{{\dagger}}[0]_{O^{\prime}_{M}(\mathcal{K}^{\prime})}$ and consider congruence (3.3) in $\mathcal{N}^{{\dagger}}[0]$:
@@
The following proposition implies Theorem 3.1 when $w=p$.
@@
Use induction on $w\geqslant 1$. If $w=1$ there is nothing to prove. Assume that proposition holds for some $1\leqslant w<p$.
For (non-commuting) variables $U$ and $V$, let
@@
Use that $\Omega\left(\sigma^{M-1}e_{1}^{(q)}\right)\circ X=\sigma X\circ\sigma^{M-1}e_{1}^{(q)}$ and there are the following congruences modulo $\left([\mathcal{L}^{(v_{0})},\mathcal{L}]+C_{w+1}(\mathcal{L})\right)_{sep}$:
@@
In order to prove part $b_{w+1})$ it will be suficient to verify that
$(\bullet_{1})$ if $p^{m}D_{a0}\in\mathcal{L}(w+1)_{W(k)}$ then $p^{m}D_{a0}\in(\mathcal{L}^{(v_{0})}+C_{w+1}(\mathcal{L}))_{W(k)}$.
This will be done by the use of the characterization of $\mathcal{L}^{(v_{0})}$ via
(3.4) $v(\mathcal{K}^{\prime}(X\,\mathrm{mod}\,\mathcal{L}^{(v_{0})}_{sep})/\mathcal{K}^{\prime})<qv_{0}-b^{*}\,,$
Suppose $y=\sum_{\iota}y_{\iota}+y_{+}$, where $y_{+}\in\mathcal{L}_{\mathrm{m}^{\prime}}$, $\iota$ runs over $\mathfrak{A}^{+}(p)$ and the corresponding $y_{\iota}$ appear in the form $y_{\iota}=\sum_{r\geqslant 0}t_{1}^{-\iota p^{r}}l_{\iota r}$ with all $l_{\iota r}\in\mathcal{L}_{W(k)}$.
Then $X=\sum_{\iota}X_{\iota}+X_{+}$ where $\sigma X_{\iota}-X_{\iota}=y_{\iota}$ and $\sigma X_{+}-X_{+}=y_{+}$. Here $X_{+}\in\mathcal{L}_{\mathrm{m}^{\prime}}$ and does not contribute to (3.4). The elements $X_{\iota}$ contribute to (3.4) as follows:
if $\iota\in\mathfrak{A}^{+}(p)$ and $\iota p^{-v_{p}(\iota)}\geqslant qv_{0}-b^{*}$ then
$\sum_{r}\sigma^{-r}l_{\iota r}\in(\mathcal{L}^{(v_{0})}+C_{w+1}(\mathcal{L}))_{W(k)}\,.$
We can follow the contribution to the upper ramification in (3.4) for each $\iota\in\mathfrak{A}^{+}(p)$ separately because by Proposiion 2.4 these contributions are independent.
Let $\iota_{0}=qp^{M-1}a-p^{M-1}b^{*}U(1,m)=p^{M-1-m}(qp^{m}a-b^{*})\in\mathfrak{A}^{0}$.
Then $\iota_{0}>0$. Indeed, $p^{m}D_{a0}\in\mathcal{L}(w+1)_{W(k)}$ implies that
$qp^{M-1}a\geqslant qp^{M-1}v_{0}U(w,m)>p^{M-1}b^{*}U(w,m)\geqslant p^{M-1}b^{*}U(1,m)\,.$
Therefore, $\iota_{0}\in\mathfrak{A}_{1}^{+}(p)$ (use that $a$ is prime to $p$) and $p^{m}a\geqslant v_{0}$. This means that the $\iota_{0}$-part $X_{\iota_{0}}$ should contribute to condition (3.4).
In order to find the contribution of $X_{\iota_{0}}$ notice that the element $y$ has a linear part
(3.5) $L(y):=\Omega\left(\sigma^{M-1}e_{1}^{(q)}\right)-\sigma^{M-1}e_{1}^{(q)}=\sum_{a,s}t_{1}^{-qp^{M-1}a}aS_{b^{*}}^{s}D_{a0}$
and a non-linear part, which is a $W(k)$-linear combination of various